h) A diode laser beam with a wavelength of 650 nm is divided into two beams and allowed to travel through two arms of a coupled-waveguide with the same length of 10 mm and with a refractive index value of 1.49. Determine the phase difference between light-beams traveling through the waveguides when the refractive index value of one of the arm changes to 1.50 with a temperature change in one of the arm. (Speed of light: c-3 x 108 m/s; Planks constant: h-6626 x104 Js)
Answer #1

Optical path difference D= difference of index* thickness=(1.50-1.49)*10=0.1 mm

Phase shift= 2*3.14*D/wavelength =2*3.14*0.1*10-3/650*10-9

=9.66*100 =966